OpenStudy (angelwings996):

Let u = <4,-1> and v = <3,5>. What is lu + vl rounded to the nearest hundredth?

OpenStudy (angelwings996):

That is suppose to be absolute values btw

OpenStudy (zzr0ck3r):

do you know the distance function?

OpenStudy (angelwings996):

distance formula?

OpenStudy (zzr0ck3r):

this is the norm of two vectors norm(a,b) = || a - b || = || b - a || = sqrt((a_1)-(b_1))^2 + (a_2 - b_2)^2 + .......)

OpenStudy (zzr0ck3r):

so you have sqrt((5-(-1))^2 + (3-4)^2) = sqrt(36+ 1) = sqrt(37)

OpenStudy (angelwings996):

Okay, is there more so do I just round that ?

OpenStudy (angelwings996):

*or not so

OpenStudy (zzr0ck3r):

37 is prime I think

OpenStudy (zzr0ck3r):

so yes round that

OpenStudy (angelwings996):

Okay, so would it be 6.1 ?

OpenStudy (zzr0ck3r):

its Pythagoreans theorem

OpenStudy (zzr0ck3r):

sure sounds about right

jimthompson5910 (jim_thompson5910):

the answer should be larger than 6.1

OpenStudy (angelwings996):

how ?

jimthompson5910 (jim_thompson5910):

u = <4,-1> and v = <3,5> u+v = <4,-1>+<3,5> u+v = <4+3,-1+5> u+v = <7,4>

OpenStudy (zzr0ck3r):

o wow I read minus, ...... sorry

jimthompson5910 (jim_thompson5910):

|u+v| = sqrt(7^2+4^2) |u+v| = sqrt(49+16) |u+v| = sqrt(65) |u+v| = 8.06225774829854 |u+v| = 8.06

OpenStudy (angelwings996):

Oh okay, that sounds a bit better. Thank you @jim_thompson5910

OpenStudy (zzr0ck3r):

change my 5- to 5+ and 3- to 3+ and your good:) sorry

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (angelwings996):

Thats okay @zzr0ck3r I get what you did though (: